# Direct-CP: Directed Collaborative Perception for Connected and Autonomous Vehicles via Proactive Attention

### v1

@@ I Introduction
Figure 1: Overview of directed CP framework. With a limited communication budget, ego CAV may hope to enhance its CP performance more in certain directions where the traffic is complex while keeping a basic perception of other directions with minimal traffic.
@@
In addition, the communication overhead is critical factor that must be carefully considered when designing a CP system[7, 8, 2, 9, 10, 11, 12, 13, 14, 15, 16]. With constraints such as a limited communication budget and a maximum allowable delay, engaging all collaborators and fully utilizing their captured views for enhancing perception across 360 degrees can significantly burden both communication and computational resources. This is particularly severe when the number of collaborators and the frame rate (measured in frames per second, FPS) are high. Indeed, reallocating communication resources from less critical directions to enhance perception in more important areas is not only strategically advantageous but also enhances CP efficiency in terms of both communication and computation.
@@ II Related Works
PACP[1] Full Priority-based average Omnidirectional
When2com[5] Full Agent-level attention Omnidirectional
@@
Where2com[10] Sparse Confidence-aware attention Omnidirectional
Direct-CP (Ours) Sparse Proactive selective attention Directed
@@ II-A Collaborative perception
Collaborative perception has gained significant attention for its ability to enhance the sensing capabilities of individual vehicles beyond the constraints of isolated sensors. This approach favors intermediate-stage fusion strategies [19, 20, 21, 22, 23], which facilitate the exchange of intermediate feature representations among CAVs to improve collaboration. However, as the dimensionality of features and the number of collaborators grow, efficient bandwidth management becomes essential. Who2com [17] introduces a sophisticated multi-stage handshake mechanism that trains neural networks to compress critical information for each stage, optimizing vehicle connectivity through a matching score. V2VNet [8] leverages a graph neural network to effectively aggregate information from nearby CAVs, significantly enhancing collective perception by establishing a robust information network. However, these approaches overlook the varying importance of individual CAVs in optimizing CP. To address this, PACP [1] implements a BEV-match mechanism to prioritize collaborative CAVs before message fusion. Nevertheless, PACP focuses solely on prioritizing each agent without considering different priority levels of various views from a single agent. Besides, PACP aims to optimize omnidirectional CP performance, lacking the flexibility to enable an ego CAV to dynamically and proactively adjust its directional CP, which is the focus of this paper.
@@ III Methodology
Figure 2: Method overview. The ego CAV integrates DAS from nearby RSU and its interest weights to create the direction mask. Subsequently, the initial query map, pose data from other CAVs, direction mask, and communication budget are fed into QC-Net to produce sparse query maps. QC-Net consists of two main components: (i) Direction Control Module to output query confidence maps (QCMs) prioritizing direction, and (ii) Query Clipping Layer to rank QCMs and generate binary query maps only selecting the top $Q_{max}\times H\times W$ significant queries yielding to communication budgets.
The overall architecture of our method is depicted in Fig. 2. Several RSUs have been deployed along the roadway to capture the traffic distribution surrounding the ego vehicle. Owing to their elevated positions, RSUs are capable of monitoring a more extensive traffic view than individual vehicles can achieve. Periodically, the ego CAV transmits its location and speed data to the nearby RSU and receives a computed direction attention score (DAS) from the RSU for reference. Utilizing DAS in conjunction with its own interests, the ego vehicle identifies which directions are temporarily non-essential and consequently masks them during the collaborative perception process. Subsequently, guided by its prioritized directions, communication budget, and the pose data of other CAVs, the ego CAV refines the selection of optimal feature map queries to nearby CAVs, aiming to maximize the directed perception performance within the constraints on communication budgets. The detailed methods are elaborated in the following subsections.
@@ III-A RSU-aided direction masking
In this paper, we resort to RSUs to aid ego CAVs in judging the important directions. We consider splitting the 360-degree space surrounding an ego CAV into $N_{dir}$ local directions. Based on the location and speed of an ego CAV, the corresponding RSU first projects it into the 2D view it captured, and then calculates DAS in $N_{dir}$ predefined directions. Here, for simplicity, we take the detected number of vehicles as the indicator to calculate DAS. Thus, the returned DASs from the RSU are represented as $\{\mathcal{S}_{r}^{i}\}_{k=1}^{N_{dir}}=\{N_{vec}^{i}\}_{k=1}^{N_{dir}}$, where $N_{vec}^{i}$ denotes the detected number of vehicles in the $i$-th direction surrounding the ego CAV. After getting the DAS $\{\mathcal{S}_{r}^{i}\}_{i=1}^{N_{dir}}$ from RSU, the ego CAV calculates the final direction mask combining its own interest weights $\{\mathcal{I}_{e}^{i}\}_{i=1}^{N_{dir}}$. The interest weights can be flexibly set according to the ego vehicle’s proactive willingness. When the interest weights are uniformly assigned, the ego vehicle determines the direction importance totally according to the RSU. The final direction mask $\{\mathcal{M}_{i}\}_{i=1}^{N_{dir}}$ is calculated as follows:
@@
where Heaviside step function $H(\cdot)$ equals 1 when parameters are positive otherwise 0. $\sigma_{1}$ is the threshold to judge if the $i$-th direction is important among all the directions. However, in some cases, there could be complex traffic even in the less important direction, so we set an additional constant threshold $\sigma_{2}$ to recognize this case. Our proposed RSU-aided direction masking mechanism provides two important insights: 1) Only the ego location, speed, and the DAS from RSU are transmitted during the interaction, requiring minimal bandwidth and keeps real-time communication; 2) The interest weight matrix fully ensures the ego vehicle’s proactivity in judging the important directions, and the ego vehicle can easily adjust interest weights to patronize the directions regardless of the RSU’s suggestions.
@@ III-C Direction-weighted detection loss
Figure 4: Visualization of Direct-CP and baselines on V2X-Sim 2.0 dataset. The green boxes are ground truth and the red boxes are predictions. Where2comm achieves higher global CP accuracy than the lower-bound but degrades in some local directions. Our proposed Direct-CP proactively guides the attention to boost perception in the ego’s interested directions (denoted with the number 1, and the right arrow indicates the ego’s moving direction).
@@ IV-B Quantitative results
Figure 5: Visualization of ego CAV’s attention weights on neighboring CAVs. Where2comm attends to features of other CAVs more equally to optimize omnidirectional perception. In contrast, our Direct-CP is direction-aware and queries features that are more relevant to the ego’s interested directions.
@@
TABLE II: Quantitative results of collaborative 3D detection (communication budget = 0.2, * indicates patronized directions).
@@
When2comm 34.97 33.56 19.36 49.96 53.56 24.81 24.59 8.75 40.42 38.70
V2VNet 57.49 53.62 28.01 60.36 67.35 42.61 41.05 19.54 36.59 48.22
Where2comm 51.29 59.38 48.83 56.27 79.59 45.86 44.52 37.89 48.83 64.96
Direct-CP 65.84 ($\uparrow 28.4\%$) 65.48 ($\uparrow 10.3\%$) 37.21 60.55 81.17 ($\uparrow 2.0\%$) 55.76 ($\uparrow 21.6\%$) 53.20 ($\uparrow 19.5\%$) 28.62 49.98 66.57 ($\uparrow 2.5\%$)
TABLE III: Ablation studies on the effect of DWLoss weight factor $\sigma$ (communication budget = 0.2).
@@ IV-C Qualitative results
In this paper, we have introduced Direct-CP, a novel CP system for ego vehicles to enhance perception in patronized directions. We have developed an RSU-aided direction masking by integrating RSU’s traffic detection with the ego vehicle’s interest weights to identify key directions. We have also designed a proactive direction-aware attention mechanism to intelligently collect sparse feature maps from multiple vehicles under limited communication budgets, thus improving local directional perception. Additionally, we have created a direction-weighted detection loss to align perception outputs with ground truth more accurately. Extensive experiments have been conducted and the results have demonstrated that Direct-CP achieves directed performance gains under constrained communication resources and outperforms baselines in terms of flexibility and efficiency.

### v2

@@ I Introduction
Figure 1: Overview of directed CP framework. Under limited communication budget, ego CAV enhances perception in complex traffic directions while maintaining basic awareness in low-traffic areas.
@@
In addition, the communication overhead is critical factor that must be carefully considered when designing a CP system[7, 8, 2, 9, 10, 11, 12, 13]. With constraints such as a limited communication budget and a maximum allowable delay, engaging all collaborators and fully utilizing their captured views for enhancing perception across 360 degrees can significantly burden both communication and computational resources. This is particularly severe when the number of collaborators and the frame rate (measured in frames per second, FPS) are high. Indeed, reallocating communication resources from less critical directions to enhance perception in more important areas is not only strategically advantageous but also enhances CP efficiency in terms of both communication and computation.
@@ II Related Works
PACP [1] Full Priority-based Omnidirectional
When2com [5] Full Agent attention Omnidirectional
@@
Where2com [10] Sparse Location attention Omnidirectional
Direct-CP (Ours) Sparse Proactive attention Directed
@@ II-A Collaborative perception
Collaborative perception extends vehicles’ sensing capabilities beyond individual sensor limits through intermediate-stage fusion strategies [16, 17, 18, 19, 20]. While this enables feature exchange among CAVs, increasing feature dimensions and collaborator numbers demand efficient bandwidth management. Who2com [14] employs a multi-stage handshake mechanism to compress information via matching scores, while V2VNet [8] uses graph neural networks to aggregate information from nearby CAVs. These methods, however, neglect the varying importance of individual CAVs. PACP [1] addresses this limitation with a BEV-match mechanism for CAV prioritization, but it only considers agent-level priorities while ignoring view-specific importance. Furthermore, PACP’s focus on omnidirectional perception lacks the flexibility for directional perception adjustment, which is the focus of this paper.
@@ III Methodology
Figure 2: Method overview. Ego CAV combines RSU’s DAS with its interests to create direction mask. The mask, along with initial query map, neighboring CAVs’ poses, and communication budget, are input to QC-Net. QC-Net contains: (i) Direction Control Module generating direction-prioritized query confidence maps (QCMs), and (ii) Query Clipping Layer selecting top $Q_{max}\times H\times W$ queries based on QCMs under budget constraints.
Fig. 2 illustrates our method’s architecture. RSUs deployed along the roadway monitor traffic from elevated positions, providing broader views than individual vehicles. The ego CAV periodically exchanges its status with nearby RSUs to receive direction attention scores (DAS). Based on DAS and its interests, the ego vehicle masks non-essential directions during collaborative perception. Then, guided by prioritized directions, communication budget, and neighboring CAVs’ poses, it selects optimal feature map queries to maximize directed perception performance. The following subsections detail these components.
@@ III-A RSU-aided direction masking
In this paper, we leverage RSUs to help ego CAVs identify important directions. The 360-degree space around an ego CAV is divided into $N_{dir}$ local directions. Based on ego CAV’s location and speed, the corresponding RSU projects it into its captured 2D view and calculates Direction Attention Scores (DAS) for each direction. For DAS calculation, we primarily use vehicle density as an indicator, as areas with higher vehicle density typically require more attention. The DAS from RSU is represented as $\{\mathcal{S}_{r}^{i}\}_{k=1}^{N_{dir}}=\{N_{vec}^{i}\}_{k=1}^{N_{dir}}$, where $N_{vec}^{i}$ denotes the number of detected vehicles in the $i$-th direction. When RSU is unavailable, the system can alternatively use historical CP results to estimate traffic density, ensuring system robustness. This approach can be extended to incorporate additional factors such as vehicle speeds, accident history, and road conditions. The ego CAV then combines RSU’s DAS $\{\mathcal{S}_{r}^{i}\}_{i=1}^{N_{dir}}$ with its own interest weights $\{\mathcal{I}_{e}^{i}\}_{i=1}^{N_{dir}}$ to calculate the final direction mask. These interest weights can be flexibly adjusted according to the ego vehicle’s proactive needs. In cases where the weights are uniformly assigned, the direction importance relies entirely on RSU’s DAS. The final direction mask $\{\mathcal{M}_{i}\}_{i=1}^{N_{dir}}$ is calculated as follows:
@@
where Heaviside step function $H(\cdot)$ equals 1 for positive inputs and 0 otherwise. The threshold $\sigma_{1}$ determines the relative importance of each direction, while $\sigma_{2}$ serves as an absolute threshold to identify complex traffic scenarios even in relatively less important directions.
Our RSU-aided direction masking mechanism offers several key advantages. First, the communication between ego vehicle and RSU only involves basic information (location, speed, and DAS), ensuring minimal bandwidth usage and real-time performance. Second, the interest weight matrix provides ego vehicles with full autonomy in direction prioritization, allowing them to override RSU suggestions when necessary. Third, the dual-threshold design ($\sigma_{1}$ and $\sigma_{2}$) enables both relative and absolute traffic complexity assessment, enhancing the system’s adaptability to various scenarios.
@@ III-C Direction-weighted detection loss
Figure 4: Visualization comparison on V2X-Sim 2.0 dataset. Green and red boxes denote ground truth and predictions, respectively. While Where2comm improves global perception over lower-bound, it shows limitations in certain directions. Direct-CP enhances perception in ego’s interested directions (marked as 1, with right arrow showing ego’s movement).
@@ IV-B Quantitative results
Figure 5: Attention weight visualization on neighboring CAVs. Where2comm distributes attention equally for omnidirectional perception, while Direct-CP focuses on features relevant to ego’s interested directions.
@@
TABLE II: Quantitative results of collaborative 3D detection (communication budget = 0.2). * indicates interested directions.
@@
When2comm [5] 34.97 33.56 19.36 49.96 53.56 24.81 24.59 8.75 40.42 38.70
V2VNet [8] 57.49 53.62 28.01 60.36 67.35 42.61 41.05 19.54 36.59 48.22
Where2comm [10] 51.29 59.38 48.83 56.27 79.59 45.86 44.52 37.89 48.83 64.96
Direct-CP (Ours) 65.84+28.4% 65.48+10.3% 37.21 60.55 81.17+2.0% 55.76+21.6% 53.20+19.5% 28.62 49.98 66.57+2.5%
TABLE III: Ablation studies on the effect of DWLoss weight factor $\sigma$ (communication budget = 0.2). * indicates interested directions.
@@ IV-C Qualitative results
In this paper, we have introduced Direct-CP, a novel CP system for ego vehicles to enhance perception in patronized directions. We have developed RSU-aided direction masking by integrating RSU’s traffic detection with ego vehicle’s interests to identify key directions. We have also designed a proactive direction-aware attention mechanism to collect sparse feature maps from multiple vehicles under limited communication budgets, improving local directional perception. Additionally, we have created a direction-weighted detection loss to align perception outputs with ground truth. Extensive experiments demonstrate that Direct-CP achieves directed performance gains under constrained communication resources and outperforms baselines in efficiency. Looking forward, our direction-aware framework opens new possibilities for adaptive perception, and future work could explore incorporating sophisticated traffic indicators and extending to extreme weather conditions and complex urban environments.